OpenStudy (anonymous):

Prove the following limit lim xsin1/x = 0 x-->0

OpenStudy (zzr0ck3r):

\[\lim_{x\rightarrow0}[x\sin(\frac{1}{x})]\]?

OpenStudy (zzr0ck3r):

-1<=sin(1/x)<=1 -x<=xsin(1/x)<=x ( or vice versa) run the limit

OpenStudy (jdoe0001):

\(\bf \lim_{x\rightarrow0} x sin\left(\frac{1}{x}\right) = 0\\ -1 < sin\left(\frac{1}{x}\right) < 1 \implies -x < xsin\left(\frac{1}{x}\right) < x\\ \color{blue}{\lim_{x\rightarrow0} x =0 \qquad \lim_{x\rightarrow0} -x =0}\\ \textit{by the squeeze theorem then }\\ \lim_{x\rightarrow0} x sin\left(\frac{1}{x}\right) = 0\)

OpenStudy (jdoe0001):

ahemm, well, or equals, one sec

OpenStudy (zzr0ck3r):

you need to show both sides so -x<=xsin(1/x)<=x and -x>=xsin(1/x)>=x we have the same limit as x approaches 0 from the left and right

OpenStudy (jdoe0001):

\(\bf \lim_{x\rightarrow0} x sin\left(\frac{1}{x}\right) = 0\\ -1 \le sin\left(\frac{1}{x}\right) \le 1 \implies -x \le xsin\left(\frac{1}{x}\right) \le x\\ \color{blue}{\lim_{x\rightarrow0} x =0 \qquad \lim_{x\rightarrow0} -x =0}\\ \textit{by the squeeze theorem then }\\ \lim_{x\rightarrow0} x sin\left(\frac{1}{x}\right) = 0\)

OpenStudy (zzr0ck3r):

now consider x negative @jdoe0001 (its trivial, but still.)

OpenStudy (jdoe0001):

right

OpenStudy (zzr0ck3r):

or by definition \[given \space\epsilon>0\\let\ \space \delta=\epsilon\\then\\|x-0|< \delta\implies\space|x*\sin(\frac{1}{x})|\le |x|< \delta = \epsilon\] thus it converges to 0 by definition of limits.

OpenStudy (zzr0ck3r):

@TheDoctorRules does this make sense?